# From Abstract to Contextual: What LLMs Still Cannot Do in Mathematics

### v1

@@ 4.1 Qualitative Analysis: Incorrect Mathematical Interpretation
To examine how such failures occur, we ask GPT‑5 to categorize the errors made by DeepSeek R1, Gemini 2.5 Pro, and Qwen3‑32B on the SG and CS sets of AIME 2024 and 2025, restricted to cases where the same models solve the original version correctly. Errors are grouped into four categories: formulation (incorrect mapping from narrative to math), calculation, logic, and other (e.g., truncation, repetition). Figure 2 shows that formulation errors dominate, accounting for roughly 80% of failures across all three models and far exceeding any other type. A typical case is given in Appendix A.5, where DeepSeek R1 fails to recognize that “the time for a gear to complete one rotation is adjustable, but it must not exceed six rotations per minute” implies the inequality $x\geq 10$, with $x$ denoting seconds per rotation. This consistent pattern confirms that formulation is the primary weakness in contextual reasoning.
@@ 5.2 Training a Dedicated Formulation Model
Figure 4: Average accuracy on ContextMATH. “w/o” denotes directly solving the scenario with the reasoning model.
@@ 6 Conclusion
Appendix A Appendix
A.1 The Use of Large Language Models (LLMs)
@@
A.2 Prompts for Data Construction
@@
This section provides the generation, verification, and revision prompts used to transform original AIME and MATH problems into Scenario Grounding (SG; see sections A.2.1, A.2.2 and A.2.3) and Complexity Scaling (CS; see sections A.2.4, A.2.5 and A.2.6) versions.
A.2.1 Scenario Grounding: Generation Prompt
@@
SG Scenario Generation Prompt Your Goal: Convert an abstract math problem into a concrete, real-world story. The new story must be mathematically identical to the original. You can draft the story first, then use the following steps to ensure accuracy and format your final output correctly. [Step 1] Map All Mathematical Components List every single mathematical element from the original problem: numbers, variables, shapes, and operations. For each, assign a specific real-world analogue. Do not omit any numbers or symbols. Bad Mapping (Vague): x $\rightarrow$ initial quantity Good Mapping (Specific): x $\rightarrow$ The initial number of barrels of oil Bad Mapping: sqrt(2) $\rightarrow$ a multiplier Good Mapping: sqrt(2) $\rightarrow$ A special efficiency boost calculated as the square root of 2 Output Format for Step 1: [Step 1] [Mathematical Element] $\rightarrow$ [Specific Real-World Analogue] [Step 2] Define the Real-World Rules of Interaction Combine the analogues from Step 1 into a coherent narrative context (e.g., engineering, physics, logistics). Explicitly define how the mathematical operations translate into actions or rules within your story. This step bridges the gap between individual components and the final narrative. Example Math: f(x) = 2x + 3 Example Rule: ”The output of a machine for any given input is found by doubling the input value and then adding a fixed calibration offset of 3.” Output Format for Step 2: [Step 2] [Mathematical Operations] $\rightarrow$ [Specific Real-World Rule]
@@
SG Scenario Generation Prompt (continued) [Step 3] Write the Final Problem Combine the elements from the previous steps into a concise story. Strict Requirements: - Perfectly Equivalent: All numbers, relationships (like perpendicularity), and constraints from the original problem must be present and unambiguous in your story. - No Math Language: Avoid mathematical terms (triangle, variable, x, cos) and symbols. Use descriptive names (a triangular field, the initial investment, a cosine-wave). - Clear Question: The final question must ask for the same value as the original (with the same output format). Output Format for Step 3: [Step 3] Question: [The complete, final real-world problem in plain text.] Now, apply this process to the following problem: {original_problem_placeholder}
A.2.2 Scenario Grounding: Verification Prompt
@@
SG Scenario Verification Prompt You are a quality checker. Your task is to verify if a real-world story is a correct and clear representation of an original math problem. Your Primary Goal: Ensure perfect mathematical equivalence in a clear and concise narrative. The story does not need to be a hyper-realistic scientific model; it only needs to be a coherent and unambiguous scenario. Please review the provided materials by considering these key questions: 1. Mathematical Integrity: Does the story perfectly preserve all numbers, relationships (e.g., perpendicularity, equality), and operations from the original problem? Is anything missing or changed? 2. Clarity & Language: Is the story easy to understand? Does it successfully avoid mathematical jargon and symbols, using clear descriptions instead? Is the final question unambiguous? 3. Conciseness: Is the story direct and to the point, without unnecessary details that overcomplicate the problem? Your Output: Please provide your review as a brief summary of your findings. After your summary, conclude with the mandatory [Overall Assessment] line. Here is the output Format: [Your findings] [Overall Assessment] - [Pass/Fail] Here is the input for your review: 1. Original Math Problem {original_problem_placeholder} 2. Conversion Mappings: 2.1 Concept Mappings: {concept_mappings_list_placeholder} 2.2 Relationship and Attribute Mappings: {relationship_mappings_list_placeholder} 3. Real-World Problem: {real_world_problem_text_placeholder}
A.2.3 Scenario Grounding: Revision Prompt
@@
A.2.4 Complexity Scaling: Generation Prompt
CS Scenario Generation Prompt Task: Enhancing Math Problem Difficulty through Contextual Embedding Your goal is to increase the difficulty of a given mathematical problem’s real-world version, while preserving its core mathematical structure and the integrity of its real-world mapping. Employ the following strategies: 1. Maintain Core Alignment Ensure all mathematical quantities, variables, and relationships directly correspond to elements within the real-world context. The underlying mathematical problem must remain solvable and equivalent to the original. 2. Embed Conditions through Layered Obfuscation Replace direct mathematical statements and values with more natural, layered, and context-rich descriptions that require deductive reasoning from the solver. * A. Contextualized Numerical Properties (Disguising Specific Values): * This technique involves disguising specific numerical values by describing them through a property they possess within a realistic, tangible scenario. The goal is for the solver to derive the original number based solely on the provided context, using common knowledge or fundamental mathematical principles. Avoid creating custom or overly obscure scenarios that would require information beyond general understanding. Apply this method only when it naturally fits the number in question; don’t force its application. * Example (for the value ’4’ hours): Rather than stating ”lasts 4 hours,” describe it as ”lasts a number of hours equivalent to the unique combinations on the drone’s two-digit pre-flight security keypad, where each digit can only be ’1’ or ’2’.” (This naturally leads to $2\times 2=4$ combinations/hours). * Example (for the value ’2’): ”The unique positive whole number that, if you square it, then subtract one time itself, and finally subtract two, the result is zero.” * Example (for the value ’16’): ”smallest integer such that C! is divisible by $2^{15}$.”
@@ A.2.4 Complexity Scaling: Generation Prompt
CS Scenario Generation Prompt (continued) * B. Relational & Functional Reverse Engineering (Disguising Variables/Formulas): * If the original problem involves analyzing a function with specific constants (e.g., f(x)=(xA)(xB)f(x)=(x-A)(x-B)...), make these constants unknown (e.g., describe them as ”labels are blurred,”, ”calibration is missing”, etc.). * Provide a minimal set of realistic data points or observed behaviors sufficient for the solver to uniquely determine these unknown constants or the full function’s form. * Example: Instead of giving $f(x)=\frac{(x-18)(x-72)...}{x}$, describe it as a system’s performance, providing $P(22)$ and $P(33)$ values, requiring deduction of the hidden constants 18 and 72. * C. Structural & Conceptual Reinterpretation (Disguising Geometric/Abstract Relations): * For geometric problems or abstract relationships, rephrase mathematical properties in terms of tangible real-world structures or processes. * Example (Symmetry): Describe a figure’s symmetry as ”a balanced design around a central axis” or ”perfectly mirrored components.” Then, we only describe half of the plot. * Example (Transformations): If there are scaling or rotation relationships, we can only describe a smaller prototype that is a scaled replica of other graphics. * Example (Equivalent Constraints): Reinterpret mathematical constraints using scenario-specific terms. For example, ”a movable point lies along the perpendicular bisector of AC” instead of $AL=CL$. Another example is translating $x_{k}+\frac{1}{x_{k}}$ to $2\cos\theta_{k}$, if we let $x_{k}=e^{i\theta_{k}}$. 3. Introduce Plausible, Irrelevant Information Incorporate details that are contextually relevant to the scenario but mathematically extraneous to the problem’s solution. This increases cognitive load and realism. Example: For a drone problem, mention its battery capacity, wing span, payload weight, or the color of the packages being delivered, provided these details do not influence the distance-speed-time calculations. 4. Language Refinement and Simplicity: Make sure all descriptions are concise and clear. Avoid redundancy and unnecessary repetition. Avoid using mathematical variables (such as x, k, A, B, etc.) directly in the situation description, but give them specific real-world meanings (such as temperature, size, threshold, etc.). Output Format: First think about how to apply these strategies to the math problem. Then output the enhanced problem in plain text after ”Enhanced Contextual Math Problem:”. Here’s the input: - Original Math Problem: {original_problem_placeholder} - Real-World Version: {real_world_problem_text_placeholder}
A.2.5 Complexity Scaling: Verification Prompt
@@
CS Scenario Verification Prompt You are tasked with verifying the quality and accuracy of a real-world version of an abstract math problem. Issue Types Defined: * EQUIVALENCE_ERROR: The enhanced problem’s mathematical core is not equivalent to the original, or it’s unsolvable. Crucially, you must solve any nested or abstract sub-problems to confirm the derived mathematical information uniquely matches the original problem’s data. * Example Suggestion: ”The definition for ’standard mission duration’ should lead to a value of 4. Consider linking it to a 2 x 2 matrix operation if the context allows.” * REALISM_ERROR: The context or embedded condition feels contrived, unnatural, or like a ”brain teaser” and isn’t a plausible real-world scenario. * Example Suggestion: ”Instead of an overly abstract or obscure derivation for a value (e.g., a specific dimension or time), consider linking it to a more common measurement, a simple calculation based on widely known facts, or a basic geometric property that naturally fits the scenario.” * CONCISENESS_ERROR: The problem statement is unnecessarily long, verbose, or unclear. * Example Suggestion: ”Condense the description of the speed increment’s derivation by directly stating the polynomial’s property more concisely.” * FORMAT_ERROR: The output doesn’t follow specified formatting (e.g., uses math variables, incorrect header, ask for different value). * Example Suggestion: ”Avoid using variable m in the context.” * OTHER: Any other issues. Review Output Format: Start your output with an ‘ASSESSMENT‘ status. If the ‘ASSESSMENT‘ is ‘FAIL‘, you must provide ‘FEEDBACK‘ in the specified parseable format. [Overall Assessment] - [Pass/Fail] If ASSESSMENT is FAIL, provide FEEDBACK like this: [FEEDBACK] - ISSUE_TYPE: [EQUIVALENCE_ERROR/REALISM_ERROR/…] - DESCRIPTION: [Concise description of the specific issue.] - LOCATION: [Optional: Specific phrase or section from the problem.] - ISSUE_TYPE: [Another ISSUE_TYPE if applicable] - DESCRIPTION: [Another concise description.] Here is the input: - Original Math Problem {original_problem_placeholder} - Real-World Problem Scenario: {real_world_problem_text_placeholder}
A.2.6 Complexity Scaling: Revision Prompt
@@
A.3 Model Evaluation Details
@@
Notably, o1-mini and GPT-5 were accessed via shared endpoints that only support a temperature of 1.0.
@@
A.4 Full Results for Open-Source Models
@@
A.5 Reasoning Error Example of DeepSeek R1
We provide an example in A.5.1 to illustrate a typical failure mode of DeepSeek R1. The model’s algebraic reasoning was largely correct, but it misinterpreted how geometric constraints in the scenario map back to the original problem, leading to an incorrect answer. This case highlights how formulation errors can propagate even when subsequent reasoning is valid.
A.5.1 Case Study: Misinterpreting Geometric Constraints
@@
The Alpha Pad is 107 meters long and 16 meters wide, and the Echo Deck is 184 meters long and 17 meters wide. The straight track is treated as the x-axis in a coordinate system, with the fountain center at (h, k) and radius r. The back corners on the fountain and the points on the track are used to determine the positions. For the Alpha Pad: - BR_A is at (0, 0). - FR_A is at (107, 0), as the distance along the length is 107 meters. - BL_A is at (0, 16), assuming the platform is above the track ($y>0$). This gives k = 8 from the circle conditions, and $r^{2}=h^{2}+64$. For the Echo Deck: - FL_E is at (b, 0). - FR_E is at (b + 17, 0), as the width is 17 meters. Incorrect mathematical interpretation - BL_E is at (b, 184), assuming the platform is above the track ($y>0$). The subsequent reasoning process is correct, but based on wrong conditions. GPT-5 analysis: Primary error type: Incorrect mathematical interpretation (with a secondary logical/rounding issue at the end). What went wrong: 1) Misassigning which side of the Echo Deck lies along the straight track. 2) Misassigning which Echo Deck side forms the chord on the circle.
An example of R1 reasoning error (continued) 3) Additional logical issue (unjustified ”nice fraction”.) … How this led to the wrong answer: - By putting 17 along the track instead of 184 and using a 17-length chord for the circle instead of 184, the computed b (the x-position of FL_E) shifted incorrectly, producing $|107-b|\approx 60.5$, rather than the correct $CE=104$. The mistake is structural (wrong mapping of sides to the line and circle), not just arithmetic.
@@
A.6 Full Results for Problem Formulation Analysis

### v2

@@ 4.1 Qualitative Analysis: Incorrect Mathematical Interpretation
To examine how such failures occur, we ask GPT‑5 to categorize the errors made by DeepSeek R1, Gemini 2.5 Pro, and Qwen3‑32B on the SG and CS sets of AIME 2024 and 2025, restricted to cases where the same models solve the original version correctly.
@@
Errors are grouped into four categories: formulation (incorrect mapping from narrative to math), calculation, logic, and other (e.g., truncation, repetition). Figure 2 shows that formulation errors dominate, accounting for roughly 80% of failures across all three models and far exceeding any other type. A typical case is given in Appendix A.4, where DeepSeek R1 fails to recognize that “the time for a gear to complete one rotation is adjustable, but it must not exceed six rotations per minute” implies the inequality $x\geq 10$, with $x$ denoting seconds per rotation. This consistent pattern confirms that formulation is the primary weakness in contextual reasoning. This presents an interesting contrast with Cheng et al. (2025): while their findings suggest that abstract formulation is not the primary bottleneck on simpler benchmarks like GSM8K (Cobbe et al., 2021a), our work demonstrates that formulation becomes a critical bottleneck once the mathematical core is embedded in complex, contextual narratives.
@@ 5.2 Training a Dedicated Formulation Model
Table 5: Average accuracy on ContextMATH. “w/o” denotes directly solving the scenario with the reasoning model.
@@ 6 Conclusion
Acknowledgement
The work described in this paper is substantially supported by a grant from the Direct Grant of Faculty of Engineering, The Chinese University of Hong Kong (Project Code: 4055275).
Appendix A Appendix
@@
The Use of Large Language Models (LLMs)
@@
A.1 Prompts for Data Construction
@@
This section provides the generation, verification, and revision prompts used to transform original AIME and MATH problems into Scenario Grounding (SG; see sections A.1.1, A.1.2 and A.1.3) and Complexity Scaling (CS; see sections A.1.4, A.1.5 and A.1.6) versions.
A.1.1 Scenario Grounding: Generation Prompt
@@
SG Scenario Generation Prompt Your Goal: Convert an abstract math problem into a concrete, real-world story. The new story must be mathematically identical to the original. You can draft the story first, then use the following steps to ensure accuracy and format your final output correctly. [Step 1] Map All Mathematical Components List every single mathematical element from the original problem: numbers, variables, shapes, and operations. For each, assign a specific real-world analogue. Do not omit any numbers or symbols. Bad Mapping (Vague): x $\rightarrow$ initial quantity Good Mapping (Specific): x $\rightarrow$ The initial number of barrels of oil Bad Mapping: sqrt(2) $\rightarrow$ a multiplier Good Mapping: sqrt(2) $\rightarrow$ A special efficiency boost calculated as the square root of 2
@@
SG Scenario Generation Prompt (continued) Output Format for Step 1: [Step 1] [Mathematical Element] $\rightarrow$ [Specific Real-World Analogue] [Step 2] Define the Real-World Rules of Interaction Combine the analogues from Step 1 into a coherent narrative context (e.g., engineering, physics, logistics). Explicitly define how the mathematical operations translate into actions or rules within your story. This step bridges the gap between individual components and the final narrative. Example Math: f(x) = 2x + 3 Example Rule: ”The output of a machine for any given input is found by doubling the input value and then adding a fixed calibration offset of 3.” Output Format for Step 2: [Step 2] [Mathematical Operations] $\rightarrow$ [Specific Real-World Rule] [Step 3] Write the Final Problem Combine the elements from the previous steps into a concise story. Strict Requirements: - Perfectly Equivalent: All numbers, relationships (like perpendicularity), and constraints from the original problem must be present and unambiguous in your story. - No Math Language: Avoid mathematical terms (triangle, variable, x, cos) and symbols. Use descriptive names (a triangular field, the initial investment, a cosine-wave). - Clear Question: The final question must ask for the same value as the original (with the same output format). Output Format for Step 3: [Step 3] Question: [The complete, final real-world problem in plain text.] Now, apply this process to the following problem: {original_problem_placeholder}
A.1.2 Scenario Grounding: Verification Prompt
@@
SG Scenario Verification Prompt You are a quality checker. Your task is to verify if a real-world story is a correct and clear representation of an original math problem. Your Primary Goal: Ensure perfect mathematical equivalence in a clear and concise narrative. The story does not need to be a hyper-realistic scientific model; it only needs to be a coherent and unambiguous scenario. Please review the provided materials by considering these key questions: 1. Mathematical Integrity: Does the story perfectly preserve all numbers, relationships (e.g., perpendicularity, equality), and operations from the original problem? Is anything missing or changed? 2. Clarity & Language: Is the story easy to understand? Does it successfully avoid mathematical jargon and symbols, using clear descriptions instead? Is the final question unambiguous? 3. Conciseness: Is the story direct and to the point, without unnecessary details that overcomplicate the problem?
SG Scenario Verification Prompt (continued) Your Output: Please provide your review as a brief summary of your findings. After your summary, conclude with the mandatory [Overall Assessment] line. Here is the output Format: [Your findings] [Overall Assessment] - [Pass/Fail] Here is the input for your review: 1. Original Math Problem {original_problem_placeholder} 2. Conversion Mappings: 2.1 Concept Mappings: {concept_mappings_list_placeholder} 2.2 Relationship and Attribute Mappings: {relationship_mappings_list_placeholder} 3. Real-World Problem: {real_world_problem_text_placeholder}
A.1.3 Scenario Grounding: Revision Prompt
@@
A.1.4 Complexity Scaling: Generation Prompt
CS Scenario Generation Prompt Task: Enhancing Math Problem Difficulty through Contextual Embedding Your goal is to increase the difficulty of a given mathematical problem’s real-world version, while preserving its core mathematical structure and the integrity of its real-world mapping. Employ the following strategies: 1. Maintain Core Alignment Ensure all mathematical quantities, variables, and relationships directly correspond to elements within the real-world context. The underlying mathematical problem must remain solvable and equivalent to the original.
@@ A.1.4 Complexity Scaling: Generation Prompt
CS Scenario Generation Prompt (continued) 2. Embed Conditions through Layered Obfuscation Replace direct mathematical statements and values with more natural, layered, and context-rich descriptions that require deductive reasoning from the solver. * A. Contextualized Numerical Properties (Disguising Specific Values): * This technique involves disguising specific numerical values by describing them through a property they possess within a realistic, tangible scenario. The goal is for the solver to derive the original number based solely on the provided context, using common knowledge or fundamental mathematical principles. Avoid creating custom or overly obscure scenarios that would require information beyond general understanding. Apply this method only when it naturally fits the number in question; don’t force its application. * Example (for the value ’4’ hours): Rather than stating ”lasts 4 hours,” describe it as ”lasts a number of hours equivalent to the unique combinations on the drone’s two-digit pre-flight security keypad, where each digit can only be ’1’ or ’2’.” (This naturally leads to $2\times 2=4$ combinations/hours). * Example (for the value ’2’): ”The unique positive whole number that, if you square it, then subtract one time itself, and finally subtract two, the result is zero.” * Example (for the value ’16’): ”smallest integer such that C! is divisible by $2^{15}$.” * B. Relational & Functional Reverse Engineering (Disguising Variables/Formulas): * If the original problem involves analyzing a function with specific constants (e.g., f(x)=(xA)(xB)f(x)=(x-A)(x-B)...), make these constants unknown (e.g., describe them as ”labels are blurred,”, ”calibration is missing”, etc.). * Provide a minimal set of realistic data points or observed behaviors sufficient for the solver to uniquely determine these unknown constants or the full function’s form. * Example: Instead of giving $f(x)=\frac{(x-18)(x-72)...}{x}$, describe it as a system’s performance, providing $P(22)$ and $P(33)$ values, requiring deduction of the hidden constants 18 and 72. * C. Structural & Conceptual Reinterpretation (Disguising Geometric/Abstract Relations): * For geometric problems or abstract relationships, rephrase mathematical properties in terms of tangible real-world structures or processes. * Example (Symmetry): Describe a figure’s symmetry as ”a balanced design around a central axis” or ”perfectly mirrored components.” Then, we only describe half of the plot. * Example (Transformations): If there are scaling or rotation relationships, we can only describe a smaller prototype that is a scaled replica of other graphics. * Example (Equivalent Constraints): Reinterpret mathematical constraints using scenario-specific terms. For example, ”a movable point lies along the perpendicular bisector of AC” instead of $AL=CL$. Another example is translating $x_{k}+\frac{1}{x_{k}}$ to $2\cos\theta_{k}$, if we let $x_{k}=e^{i\theta_{k}}$. 3. Introduce Plausible, Irrelevant Information Incorporate details that are contextually relevant to the scenario but mathematically extraneous to the problem’s solution. This increases cognitive load and realism. Example: For a drone problem, mention its battery capacity, wing span, payload weight, or the color of the packages being delivered, provided these details do not influence the distance-speed-time calculations. 4. Language Refinement and Simplicity: Make sure all descriptions are concise and clear. Avoid redundancy and unnecessary repetition.
CS Scenario Generation Prompt (continued) Avoid using mathematical variables (such as x, k, A, B, etc.) directly in the situation description, but give them specific real-world meanings (such as temperature, size, threshold, etc.). — Output Format: First think about how to apply these strategies to the math problem. Then output the enhanced problem in plain text after ”Enhanced Contextual Math Problem:”. Here’s the input: - Original Math Problem: {original_problem_placeholder} - Real-World Version: {real_world_problem_text_placeholder}
A.1.5 Complexity Scaling: Verification Prompt
@@
CS Scenario Verification Prompt You are tasked with verifying the quality and accuracy of a real-world version of an abstract math problem. Issue Types Defined: * EQUIVALENCE_ERROR: The enhanced problem’s mathematical core is not equivalent to the original, or it’s unsolvable. Crucially, you must solve any nested or abstract sub-problems to confirm the derived mathematical information uniquely matches the original problem’s data. * Example Suggestion: ”The definition for ’standard mission duration’ should lead to a value of 4. Consider linking it to a 2 x 2 matrix operation if the context allows.” * REALISM_ERROR: The context or embedded condition feels contrived, unnatural, or like a ”brain teaser” and isn’t a plausible real-world scenario. * Example Suggestion: ”Instead of an overly abstract or obscure derivation for a value (e.g., a specific dimension or time), consider linking it to a more common measurement, a simple calculation based on widely known facts, or a basic geometric property that naturally fits the scenario.” * CONCISENESS_ERROR: The problem statement is unnecessarily long, verbose, or unclear. * Example Suggestion: ”Condense the description of the speed increment’s derivation by directly stating the polynomial’s property more concisely.” * FORMAT_ERROR: The output doesn’t follow specified formatting (e.g., uses math variables, incorrect header, ask for different value). * Example Suggestion: ”Avoid using variable m in the context.” * OTHER: Any other issues. Review Output Format: Start your output with an ‘ASSESSMENT‘ status. If the ‘ASSESSMENT‘ is ‘FAIL‘, you must provide ‘FEEDBACK‘ in the specified parseable format. [Overall Assessment] - [Pass/Fail]
CS Scenario Verification Prompt (continued) If ASSESSMENT is FAIL, provide FEEDBACK like this: [FEEDBACK] - ISSUE_TYPE: [EQUIVALENCE_ERROR/REALISM_ERROR/…] - DESCRIPTION: [Concise description of the specific issue.] - LOCATION: [Optional: Specific phrase or section from the problem.] - ISSUE_TYPE: [Another ISSUE_TYPE if applicable] - DESCRIPTION: [Another concise description.] Here is the input: - Original Math Problem {original_problem_placeholder} - Real-World Problem Scenario: {real_world_problem_text_placeholder}
A.1.6 Complexity Scaling: Revision Prompt
@@
A.2 Model Evaluation Details
@@
o1-mini and GPT-5 were accessed via shared endpoints that only support a temperature of 1.0.
@@
A.3 Full Results for Open-Source Models
@@
A.4 Reasoning Error Example of DeepSeek R1
We provide an example in A.4.1 to illustrate a typical failure mode of DeepSeek R1. The model’s algebraic reasoning was largely correct, but it misinterpreted how geometric constraints in the scenario map back to the original problem, leading to an incorrect answer. This case highlights how formulation errors can propagate even when subsequent reasoning is valid.
A.4.1 Case Study: Misinterpreting Geometric Constraints
@@
An example of R1 reasoning error (continued) The Alpha Pad is 107 meters long and 16 meters wide, and the Echo Deck is 184 meters long and 17 meters wide. The straight track is treated as the x-axis in a coordinate system, with the fountain center at (h, k) and radius r. The back corners on the fountain and the points on the track are used to determine the positions. For the Alpha Pad: - BR_A is at (0, 0). - FR_A is at (107, 0), as the distance along the length is 107 meters. - BL_A is at (0, 16), assuming the platform is above the track ($y>0$). This gives k = 8 from the circle conditions, and $r^{2}=h^{2}+64$. For the Echo Deck: - FL_E is at (b, 0). - FR_E is at (b + 17, 0), as the width is 17 meters. Incorrect mathematical interpretation - BL_E is at (b, 184), assuming the platform is above the track ($y>0$). The subsequent reasoning process is correct, but based on wrong conditions. GPT-5 analysis: Primary error type: Incorrect mathematical interpretation (with a secondary logical/rounding issue at the end). What went wrong: 1) Misassigning which side of the Echo Deck lies along the straight track. 2) Misassigning which Echo Deck side forms the chord on the circle. 3) Additional logical issue (unjustified ”nice fraction”.) … How this led to the wrong answer: - By putting 17 along the track instead of 184 and using a 17-length chord for the circle instead of 184, the computed b (the x-position of FL_E) shifted incorrectly, producing $|107-b|\approx 60.5$, rather than the correct $CE=104$. The mistake is structural (wrong mapping of sides to the line and circle), not just arithmetic.
@@
A.5 Full Results for Problem Formulation Analysis